for the pair of functions, find te indicated sum, difference, product, or quotient
f(x)=x+5,g(x)=x-4
find(f+g)(-4)

Answer by jim_thompson5910(35256)


First add the functions

f(x)+g(x)=(x+5)+(x-4)
f(x)+g(x)=2x+1
(f+g)(x)=2x+1

Then plug in x = -4

(f+g)(x)=2x+1
(f+g)(-4)=2(-4)+1
(f+g)(-4)=-8+1
(f+g)(-4)=-7